The McGraw Company is accumulating data to be used in preparing its annual profit plan for the coming year. The cost behavior pattern of the maintenance costs must be determined. The accounting staff has suggested that linear regression be employed to derive an equation in the form of y = a + bx for maintenance costs. Data regarding the maintenance hours and costs for last year and the results of the regression analysis are as follows: (CMA adapted)

Hours of Activity Maintenance Costs
January 580 $ 4,260
February 350 3,060
March 460 3,660
April 330 2,880
May 560 4,410
June 354 3,020
July 350 3,090
August 580 4,530
September 550 4,320
October 530 4,110
November 400 3,360
December 380 3,220
Sum 5,424 $ 43,920
Average 452 3,660
A coefficient 895.24
B coefficient 6.1167
Standard error of the a coefficient 108.488
Standard error of the b coefficient 0.23465
Standard error of the estimate 79.009
R2 0.98550
T-value a 8.252
T-value b 26.067


What is the fixed cost per month using the high-low method to estimate the cost equation?

Multiple Choice

  • $4,180.

  • $1,110.

  • $2,110.

  • $1,058.
Answer #1
Under High low method,Consider the Highest and Lowest activity and then apply the following equation
Variable cost per hour=Change in Total maintenance costs/Change in hours
Fixed cost=Total maintenance costs-Total variable cost
Highest activity=580 hours in January at $ 4260
(In August also 580 hours used.But in order to arrive at the answer given in the option we have to select Jan)
Lowest activity=330 hours in April at $ 2880
Variable cost per machine hour=(4260-2880)/(580-330)=1380/250=$ 5.52 per hour
To find fixed cost, Apply variable cost per unit in Lowest activity or highest activity
Let's consider Lowest activity
Fixed cost=2880-(330*5.52)=2880-1821.6=1058.4=$ 1058
Answer is 4th option. $ 1058